A regular milk that is 10% fat mix with non-fat milk to obtain a lowfat milk that is 3%. How many pounds of regular milk and how many pound of non-fat milk should be mixed to obtain 80 pounds of lowfat milk?
Answer by ptaylor(2198)

Let x=amount of regular 10% milk needed
Then 80-x=amount of non-fat milk needed
Now we know that the amount of fat in the regular milk (0.10x) plus the amount of fat in the nonfat milk (0) has to equal the amount of fat in the final mixture (0.03*80). So our equation to solve is:
0.10x=0.03*80 or
0.10x=2.4
x=24 pounds------amount of regular milk needed
80-x=80-24=56 pounds ----amount of non-fat milk needed
CK
24*0.10=0.03*80
2.4=2.4
Hope this helps---ptaylor